Flesch Corporation produces and sells two products. In the most recent month, Product C90B had sales of $24,700 and variable expenses of $6,175. Product Y45E had sales of $35,280 and variable expenses of $15,876. The fixed expenses of the entire company were $17,300. If the sales mix were to shift toward Product C90B with total dollar sales remaining constant, the overall break-even point for the entire company:

Multiple Choice

  • would increase.

  • would not change.

  • could increase or decrease.

  • would decrease.
Answer #1

Contribution margin ratio = Contribution margin / Sales

Contribution margin ratio

C90B = (24,700-6,175)/24,700 = 75%

Y45E = (35,280-15,876)/35,280 = 55%

Higher the contribution margin ratio, lower the Breakeven point

So, if sales mix shifts to product C90B, overall Breakeven point Decreases

Option D